OpenStudy (anonymous):

The square root of 114 lies between which 2 numbers?

OpenStudy (unklerhaukus):

i know that \[10^2=100\] i also know that \[11^2=121\]

OpenStudy (unklerhaukus):

another way to right this information is \[\sqrt {100}=10,\quad\sqrt{121}=11\]

OpenStudy (unklerhaukus):

\[\sqrt{100}<\sqrt{114}<\sqrt{121}\]